6. Given the following, what is the after-tax cash flow? Assume No Cap. Ex and no principal payments

Cost of goods 100

Depreciation and Amortization 35

Revenues 150

Selling, General, and Admin Exp 5

Tax rate 30%

After cash flow =

Use this data for the next two problems

Sales $100.00

COGS 0.2

General and Admin $15.00

Depreciation $35.00

Interest Expense $35.00

Tax rate 0.25

7. Is the company profitable and by how much? (yes/ no)

8. Does the company produce a positive cash flow or negative? (yes or now and how much?)

11. You are a venture capitalist with investments in 10 distinct and separate entities. Viewing your portfolio in total, if your portfolio goes up 50% this year and down 50% next year, if you started with $100, what value would you have at the end of next year?

12. Rank the following according to Payback (assume cash flows are uniform throughout the period)

“N” $(5,000,000) $350,000 $700,000 $1,500,000 $4,000,000

NPV for Project “N” = $238,013 IRR for Project “N” 16.4%

“O” $(1,500,000) $600,000 $300,000 $650,000 $900,000 $450,000

NPV for Project “O” = $414,275 IRR for Project “O” 25.6%

“P” $(4,000,000) $1,500,000 $1,875,000 $2,500,000 $350,000 $(250,000)

NPV for Project “P” = $441,727 IRR for Project “P” 21.0%

“Q” $(2,000,000) $1,580,000 $600,000 $45,000 $45,000 $45,000

NPV for Project “Q” = $(94,711) IRR for Project “Q” 11.0%

13. Value of House = 220,000

      Mortgage= 200,000

      Rate= 6%

      Term years = 30 FV House

                              $1,199.10 Mortgage Balance

Inflation on the house value is 3% per year (Compound yearly)

What is the equity in the house (value less mortgage outstanding) at the end of year 5?

14. Scrap metal has a market value of $80 per ton

A 25-year-old oil tanker has 25,000 tons of metal

It will cost you $125,000 to deliver the tanker to the scrap dealer

The tanker is still usable and generates $155,000 after tax cash

It is assumed the tanker can be used for another 10 years (worthless on the last day 10 yrs from now)

You are offered $1,750,000 for the tanker (today)

Which valuation method has the most value; assuming a 15% return required?

15. You start an engineering consulting firm. At the end of the first year, the cash flow of the firm is $75,000. The cash flow (net of all costs and expenses) grows at a rate of 15% per year. In year 7, a national engineering firm offers to buy your local firm for 5 times cash flow. What is the price the national firm will pay for your company?
Answer #1

6. The following values have been provided in the question:

Revenues = 150

Cost of goods = 100

Depreciation and Amortization = 35

Selling, General, and Admin Exp = 5

Tax rate = 30%

Solution:

After tax cash flow = Net profits after tax + Non cash expenses

Revenues = 150

Less: Cost of goods = (100)

Less: Depreciation and amortisation = (35)

Less: Selling, general and Admin exp = (05)

Profits before taxes = 10

Less: Taxes [10*30%]    = (03)

Profits after taxes = 07

Add: Depreciation & Amortization [Note 1] = 35

Cash flows after taxes    = 42

Therefore, the after tax cash flows are 42.

Note: Depreciation and Amortization expenses have been added back since they are non cash expenses.

7. The following values have been provided in the question:

Sales $100.00

COGS 0.2 (assuming this as a percentage of sales)

General and Admin $15.00

Depreciation $35.00

Interest Expense $35.00

Tax rate 0.25

Solution:

Profits before taxes = Sales - COGS - General and Admin - Depreciation - Interest expenses

= 100 - (0.2*100) - 15 - 35 - 35

= -5

In this case, since the company is incurring losses there is no need to pay taxes. Hence the profits after taxes will also be -5.

Since the profits after taxes is -5, the company is not a profitable one.

8. Cash flow = Profits after taxes (as computed in 7 above) + Non cash expenses (i.e. Depreciation)

= -5 +35

= 30

Therefore, the company is making a positive cash flow of $30.

11. Details provided in the question:

Value of portfolio at the beginning of year 1 = $100

Increase in value of portfolio during year 1 = 50%

Decrease in the value of portfolio during year 2 = 50%

Solution:

Portfolio value at the beginning of year 1 [A] = $100

Add: Increase in portfolio value during year 1 [ $100*50% ] [B] = $50

Portfolio value at the end of year 1 [C=A+B] = $150

Less: Decrease in portfolio value during year 2 [ $150*50% ] [D] = $75

Portfolio value at the end of year 2 [E= C-D] = $75

Therefore, value of portfolio at the end of year 2 is $75

12. Pay back period = E+

Where 'E' = Number of years immediately before the recovery year

'B'= Balance amount to be recovered

'C' = Cash inflow in the year of recovery

Project N = 5000000 (350000+700000+ 1500000) 3years 4000000

= 3 years + 0.6125 years

= 3.6125 years

Project O = 1500000 (600000+300000) 2years 650000

= 2.92 years

Project P = [4000000 (1500000+1875000) 2years 2500000

= 2.25 years

Project Q = [2000000 (1580000)] 1 years 600000

= 1.7 years

The project having the least payback period would be preferred first. Therefore, the ranking would be as follows:

Ranks Project PBP
1 Q 1.7 years
2 P 2.25 years
3 O 2.92 years
4 N 3.6125 years